OpenStudy (alyssa_xo):

Flip a coin until heads shows, assume that the probability of heads on one flip is \(\large \frac{3}{5}\). We define a RV (random variable) X = the # of flips. 1. What are the possible values of X? 2. Find the probability distribution function for X: give the first four values then find a general formula for the probability that X = n.

OpenStudy (dumbcow):

this is a geometric distribution http://en.wikipedia.org/wiki/Geometric_distribution X = 1,2,3,4 ... \[P(X=n) = (\frac{3}{5}) (\frac{2}{5})^{n-1}\]

OpenStudy (alyssa_xo):

So "possible values of X" just wants a generic formula? Thank you. If I wanted to sum this up using the formula for the sum of a geometric series, would this work? \(\huge \frac{3}{5}*\frac{1-\frac{2}{5}^n}{1-\frac{2}{5}}\)?

OpenStudy (dumbcow):

yes that would work :)

OpenStudy (alyssa_xo):

n or n-1?

OpenStudy (dumbcow):

you can simplify that though .... 1-2/5 = 3/5 the 3/5 cancel

OpenStudy (dumbcow):

n

OpenStudy (alyssa_xo):

ah, so we're left with \(\huge \frac{3}{5}\frac{\frac{3}{5}^n}{\frac{3}{5}}=\frac{3}{5}*\frac{3}{5}^{n-1}\)?

OpenStudy (alyssa_xo):

that's probably wrong because of pemdas bleh

OpenStudy (dumbcow):

oh careful \[1- a^n \ne (1-a)^n\]

OpenStudy (dumbcow):

also if you go to wiki link the sum should equal the cumulative distribution ----> 1 - (2/5)^n

OpenStudy (alyssa_xo):

We have to prove that the sum of all the probabilities is 1. \(\large \cancel{\frac{3}{5}}*\frac{1-\frac{3}{5}^n}{\cancel{\frac{3}{5}}}\)

OpenStudy (alyssa_xo):

\(\large \cancel{\frac{3}{5}}*\frac{1-\frac{2}{5}^n}{\cancel{\frac{3}{5}}}\) fixed

OpenStudy (dumbcow):

oh ok thats easy, the sum of all probabilities is the infinite sum of the series let n ->infinity , what happens?

OpenStudy (alyssa_xo):

1-0? oh wow, thanks

OpenStudy (dumbcow):

yep, yw :)